mary has a total of 30 dimes and nickels in her pocket worth $2.35. How many of each coin does she have?

i first had n represent the number of nickles and d number of dimes so i did n+d=30 but now im not to sure of what to do next?

Answer by Alan3354(69443)

mary has a total of 30 dimes and nickels in her pocket worth $2.35. How many of each coin does she have?
i first had n represent the number of nickles and d number of dimes so i did n+d=30 but now im not to sure of what to do next?
--------------
Make a 2nd eqn for the amounts, or value
5n + 10d = 235 --> n + 2d = 47
Can you do the rest?